OpenStudy (anonymous):

find a formula for f'(x) and determine the slope f'(a) at the point where x=a is given. f(x)= 5*e^x/2 ; x=0

OpenStudy (e.mccormick):

"slope f'(a) at" so the slope of the derivative? Or is that supposed to be "slope f(a) at"?

OpenStudy (e.mccormick):

Ah, ok... I added an OF there mentally.

OpenStudy (e.mccormick):

OK, \(f'(x)\) will give you the slope, so what you really need is \(f'(x)\). It is the answer to the first part and finds the second.

OpenStudy (e.mccormick):

So, what is the derivative of \(\Large5e^{\frac{x}{2}}\)? Do you know how to find that one.

OpenStudy (anonymous):

I have no idea because of the e^x/2. what formula would I use?

OpenStudy (e.mccormick):

OK, so, do you know the derivative of just \(e^x\)?

OpenStudy (anonymous):

no

OpenStudy (e.mccormick):

OK. That one is easy. The derivative of \(e^x\) is \(e^x\). It is special. It is its own derivative.

OpenStudy (anonymous):

so how would I find e^x/2

OpenStudy (e.mccormick):

This will be a chain rule. You have \(f(x)^{g(x)}\). Now the \(f(x)\) part is \(5e^{x/2}\) and the \(g(x)\) part is \(\frac{x}{2}\).

OpenStudy (e.mccormick):

The chain rule is for compositions of functions. Because \(e^x\) is a function and the power is made of another function, you do the chain rule. Do you remember the chain rule?

OpenStudy (anonymous):

is it e^(g(x)) * g '(x) ?

OpenStudy (e.mccormick):

!! Exactly!

OpenStudy (anonymous):

but how do I find g '(x) if g(x)= x/2?

OpenStudy (e.mccormick):

Well, it is a variable over a constant, but there is another way to think about that. When you multiply x by a fraction, what happens to it? And can you undo the multiplication?

OpenStudy (anonymous):

im confused...

OpenStudy (e.mccormick):

\[5\cdot \frac{1}{3}=\frac{5}{3}\]So what if that 5 was your x and that 3 was your 2?

OpenStudy (anonymous):

x*1/2?

OpenStudy (e.mccormick):

Yep! It just hopps off the /2 and makes a half times x.

OpenStudy (anonymous):

so would the equation be 5e^x/2 * 1/2x ?

OpenStudy (e.mccormick):

Almost, you need to differentiate the \(\frac{1}{2}x\)

OpenStudy (anonymous):

so 1/2

OpenStudy (e.mccormick):

/cheer Yep.

OpenStudy (e.mccormick):

And since the differentiatoion of the e part is itself, it is just the original equation, but now over 2!

OpenStudy (anonymous):

so f '(x)=(5e^x/2) / 2

OpenStudy (e.mccormick):

Yes. That is the first half of this. When it comes to differentiation and integration, e in \(e^x\) stands for easy! It is itself. If there is anything other than just an x in the exponent, you chain rule it. And if you somehow get \(e^c\) where c is any constant, that is just a number so the rules for constants apply.

OpenStudy (e.mccormick):

So now they want you to put 0 in for x.

OpenStudy (anonymous):

would that cancel out the e since its zero?

OpenStudy (e.mccormick):

That is a good way to say it. It makes value 1.

OpenStudy (e.mccormick):

\[\frac{5e^{\frac{x}{2}}}{2}\implies \frac{5e^{\frac{0}{2}}}{2}\implies \frac{5e^0}{2}\implies \frac{5\cdot 1}{2}\implies \frac{5}{2}\]

OpenStudy (e.mccormick):

Which is exactly what you said... just all written out in math stuff. Hehe.

OpenStudy (anonymous):

ok I think I have it now thanx for your patience

OpenStudy (e.mccormick):

No problem. You just needed a little help with e and a reminder on the meaning of a fraction. Work with some more e ones and you will remember that part. The fraction... well... calculus uses every piece of math before it. Want to do good in calculus? Practice everything that came before it!

OpenStudy (e.mccormick):

I am taking an intro to Linear Algebra class at the moment... final on Monday... and it is tons and tons of basic algebra all wrapped around a few new concepts. Where everyone has problems is with dropping a sign, doing the algebra wrong, and so on.